Calculate the final specific heat capacity (SHC) of the mixture of 30g Nacl (c=867 j/kgk) and of 30g Nacl(c=867j/kgk) and 5g of kcl (c=682 j/kgk).
1
Expert's answer
2020-12-02T09:44:36-0500

"c_{mix}=c_1\frac{m_1}{m}+c_2\frac{m_2}{m}=867\cdot\frac{0.03}{0.035}+682\cdot\frac{0.005}{0.035}=841(J/kg\cdot K)"